In a biology class, a group of students brought 9 mice for their experiment. They measured the weight and the length of the body of each mouse from the tail to the nose. The findings are recorded below:


Length in cm

4

5

5

5

7

8

9

9

10

Weight in kg

6

7

8

10

10

11

12

12

13


Calculate the Pearson correlation coefficient. Round your answers to the nearest hundredths.


1
Expert's answer
2022-05-31T11:13:40-0400

In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 4 & 6 & 24 & 16 & 36 \\
 \hdashline
 & 5 & 7 & 35 & 25 & 49 \\
 \hdashline
 & 5 & 8 & 40 & 25 & 64 \\
 \hdashline
 & 5 & 10 & 50 & 25 & 100 \\
 \hdashline
 & 7 & 10 & 70 & 49 & 100 \\
 \hdashline
 & 8 & 11 & 88 & 64 & 121 \\
 \hdashline
 & 9 & 12 & 108 & 81 & 144 \\
 \hdashline
 & 9 & 12 & 108 & 81 & 144 \\
 \hdashline
 & 10 & 13 & 130 & 100 & 169 \\
 \hdashline
Sum= & 62 & 89 & 653 & 466 & 927 \\
 \hdashline
\end{array}""\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{62}{9}=6.89"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{89}{9}=9.89"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=466-\dfrac{62^2}{9}=38.89"




"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=927-\dfrac{89^2}{9}=46.89"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=653-\dfrac{62(89)}{9}=39.89"




"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{653-\dfrac{62(89)}{9}}{\sqrt{(466-\dfrac{62^2}{9})(927-\dfrac{89^2}{9})}}"

"=0.93"